5A AND 5B

Feather Friends, Incorporated, distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

 

Sales$ 2,000,000
Variable expenses1,000,000
Contribution margin1,000,000
Fixed expenses180,000
Net operating income$ 820,000

 

Required:

Answer each question independently based on the original data:

 

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. Assume this year’s unit sales and total sales increase by 44,000 units and $3,520,000, respectively. If the fixed expenses do not change, how much will net operating income increase?

 

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's unit sales to increase by 10%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

 

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $60,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

 

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's unit sales by 25%. How much could the president increase this year's advertising expense and still earn the same $820,000 net operating income as last year?
Answer #1

Let's answer each question based on the given data:


1. What is the product's CM ratio? CM ratio (Contribution Margin ratio) = (Contribution Margin / Sales) * 100 CM ratio = ($1,000,000 / $2,000,000) * 100 CM ratio = 50%


2. Use the CM ratio to determine the break-even point in dollar sales. Break-even point (in dollar sales) = Fixed expenses / CM ratio Break-even point = $180,000 / 0.50 Break-even point = $360,000


3. Assume this year’s unit sales and total sales increase by 44,000 units and $3,520,000, respectively. If the fixed expenses do not change, how much will net operating income increase? Increase in unit sales = 44,000 units Increase in total sales = $3,520,000

Contribution margin per unit = Contribution Margin / Unit sales Contribution margin per unit = $1,000,000 / 44,000 units Contribution margin per unit = $22.73

Increase in net operating income = (Contribution margin per unit * Increase in unit sales) - Increase in fixed expenses Increase in net operating income = ($22.73 * 44,000) - $0 (since fixed expenses do not change) Increase in net operating income = $1,000,120


4-a. What is the degree of operating leverage based on last year's sales? Degree of Operating Leverage = Contribution Margin / Net Operating Income Degree of Operating Leverage = $1,000,000 / $820,000 Degree of Operating Leverage ≈ 1.22


4-b. Assume the president expects this year's unit sales to increase by 10%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year? Percentage increase in net operating income = Degree of Operating Leverage * Percentage increase in unit sales Percentage increase in net operating income = 1.22 * 10% Percentage increase in net operating income ≈ 12.2%


5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $60,000 increase in advertising, would increase this year's unit sales by 25%.


    a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented? New selling price = $80 - 13% of $80 = $80 - $10.40 = $69.60 New unit sales = 25% increase = 44,000 units * 1.25 = 55,000 units

    Total sales with new price and increased unit sales = $69.60 * 55,000 units = $3,828,000 New Contribution Margin = Total sales - Variable expenses = $3,828,000 - $1,000,000 = $2,828,000

    New net operating income = New Contribution Margin - Fixed expenses - New Advertising expense New net operating income = $2,828,000 - $180,000 - $60,000 = $2,588,000

    

    b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year? Net operating income increase = New net operating income - Last year's net operating income Net operating income increase = $2,588,000 - $820,000 = $1,768,000


6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's unit sales by 25%. How much could the president increase this year's advertising expense and still earn the same $820,000 net operating income as last year?

Let's first calculate the contribution margin per unit with the increased sales commission:

New contribution margin per unit = Contribution Margin per unit - Increase in sales commission New contribution margin per unit = $22.73 - $1.80 = $20.93

Now, we can calculate the required increase in unit sales to achieve the same net operating income as last year:

Required increase in unit sales = (Fixed expenses + Advertising expense) / New contribution margin per unit Required increase in unit sales = ($180,000 + Advertising expense) / $20.93

Since the increase in unit sales is 25%, we can write:

1.25 * Current unit sales = ($180,000 + Advertising expense) / $20.93

Now, solve for Advertising expense:

Advertising expense = 1.25 * Current unit sales * $20.93 - $180,000

Use the given information of last year's unit sales (let's assume it's X units):

Advertising expense = 1.25 * X * $20.93 - $180,000

Set this equal to the last year's net operating income to find the value of X:

1.25 * X * $20.93 - $180,000 = $820,000

1.25 * X * $20.93 = $820,000 + $180,000

1.25 * X * $20.93 = $1,000,000

X ≈ $1,000,000 / (1.25 * $20.93)

X ≈ $1,000,000 / $26.1625

X ≈ 38,199 units (approx.)


So, the president could increase this year's advertising expense to achieve the same $820,000 net operating income as last year if the unit sales increase to approximately 38,199 units.

 


answered by: Aratrika